OpenStudy (lolipop_drum312):

Ok new mothersucking question.... help please :) Convert 50 ft per s to mph. 5280 ft = 1 mi Round to the nearest hundredth. Enter your answer in the box.

OpenStudy (vheah):

This isn't too bad. Think of units you can cancel. The way I convert my units is through a method called "dimensional analysis" as my teachers would call it, or the "train track method", or whatever. So this is what you have (50 feet per s) \[\frac{ 50 ft}{ s }\] as In for every 50 feet, you have 1 second. Now, mph in essence means > mile(s) per hour. \[\frac{ mile}{ hour }\] We know that in 5280 ft = 1 mile So we mak a ratio out of the 5280 = 1 mile, to where we can cancel the unit ft on our 50 feet per s \[\frac{ 50 ft }{ s } * \frac{ 1 mile }{ 5280 ft }\] The feet cancels out and we are left with \[\frac{ 50 * 1 mile }{ 5280*s }\] = \[\frac{ 50 miles }{ 5280 s }\] if we simply that we get: 0.00947 miles per s Now we are not done yet because h (hours) in the denominator and not s. right now we still have s in the denominator. We know that in 1 hour = 60 min and in 1 min = 60 seconds So you do the same thing you did above, except we have to do it twice: \[\frac{ 0.00947 miles }{s } * \frac{ 60s }{ 1 minute } * \frac{ 60 minutes }{ 1 hour }\] the seconds cancel out since you have an 's' in the numerator and the denominator, minutes also cancel out for the same reason, and you're left with 1 hour. Multiply out and you get \[\frac{ 34.092 miles}{ hour } \] or or 34.092mph

OpenStudy (lolipop_drum312):

ooh thank you i will write this down

OpenStudy (vheah):

You're welcome :)